### v2

# Coverage errors for Student’s $t$ confidence intervals comparable to those in Hall (1988)

Abstract
Table 1 of Hall, (1988) contains asymptotic coverage error formulas for some nonparametric approximate 95% confidence intervals for the mean based on $n$ IID samples. The table includes an entry for an interval based on the central limit theorem using Gaussian quantiles and the Gaussian maximum likelihood variance estimate. It is missing an entry for the very widely used Student’s $t$ confidence intervals. This note develops such a formula. The impetus to revisit this issue arose from the surprisingly robust performance of confidence intervals based on Student’s t statistic in randomized quasi-Monte Carlo sampling. Hall’s table had $3.350.14\kappa-2.12\gamma^{2}-$ 3.35 for normal theory intervals; the corresponding entry for Student’s $t$ is $0.14\kappa-2.12\gamma^{2}$.
An earlier version of this note reported that it corrected some coverage error formulas in Hall, (1988). Two-sided errors take the form $2\Phi^{-1}(0.975)(A\kappa+B\gamma^{2}+C)\varphi(1.96)/n+O(1/n^{3/2})$ where the error may well be $O(n^{-2})$. Hall’s table showed $\Phi^{-1}(0.975)(A\kappa+B\gamma^{2}+C)$. The version intended as a correction had $2(A\kappa+B\gamma^{2}+C)$, wider by about $2/1.96\doteq 1.02$. So, Hall’s table really is proportional to the two-sided coverage errors.
1 Introduction
Table 1 of Hall, (1988) gives expressions for the coverage error in several nonparametric confidence interval methods for the mean of a real-valued random variable $X$ based on an IID sample of $n$ realizations of that random variable. It is missing an entry for the usual confidence interval based on Student’s $t$ statistic. It does have an entry for the case where the $t$ threshold is replaced by one from the standard Gaussian distribution and the usual unbiased sample variance is replaced by the maximum likelihood estimate (MLE). This note uses an expression from page 949 of that paper to develop a formula for the Student $t$ intervals.
The motivation to revisit this problem comes from recent work in randomized quasi-Monte Carlo (RQMC) sampling in L’Ecuyer et al., (2023). RQMC provides an unbiased estimate of a multidimensional integral. The sample values that go into that estimate have a complex dependence structure that improves the convergence rate beyond that of plain Monte Carlo. Then a small number of independent replicates are made in order to estimate error. The surprise in L’Ecuyer et al., (2023) was that a plain Student’s $t$ confidence interval proved to be remarkably robust in terms of coverage accuracy at 95% nominal coverage. It performed better than two bootstrap confidence intervals. A theoretical comparison of bootstrap confidence intervals from Hall, (1988) revolves around the skewness and kurtosis of the random variable being replicated. This paper fills in the missing formula for Student’s $t$ distribution.
For $n\geqslant 2$, let $X_{1},\dots,X_{n}$ be IID random variables with the same distribution as $X$. Then for $20<\alpha<1/$ 2, a standard Student’s $t$ based confidence interval for $\mu$ with nominal coverage $1-2\alpha$ takes the form
$\bar{X}\pm t^{1-\alpha}_{(n-1)}s/\sqrt{n}$ (1)
where
$\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}\quad\text{and}\quad s^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
and $t^{1-\alpha}_{(n-1)}$ is the $1-\alpha$ quantile of Student’s $t$ distribution on $n-1$ degrees of freedom. This interval has asymptotic coverage probability $1-2\alpha$. It is more common to write about coverage $1-\alpha$ using $t_{(n-1)}^{1-\alpha/2}$ for $\alpha\in(0,1)$ but the usage above conforms to Table 1 of Hall, (1988).
An alternative interval is
$\bar{X}\pm\Phi^{-1}(1-\alpha)\hat{\sigma}/\sqrt{n}$ (2)
where $\Phi$ is the cumulative distribution function of the $\mathcal{N}(0,1)$ distribution, and
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})^{2}$
is the MLE of $\sigma^{2}$ in the model $X_{i}\sim\mathcal{N}(\mu,\sigma^{2})$. This is the interval that appears in Table 1 of Hall, (1988) with the label ‘Norm’.
Table 1 of Hall, (1988) includes asymptotic coverage error formulas for ‘Norm’ and some bootstrap confidence intervals. For the necessary assumptions on the distribution of $X$, see Hall, (1988). These table entries are for $\alpha=0.025$ corresponding to the common default of 95% confidence. The coverage error of (2) for $\alpha=0.025$ is
$\frac{2}{n}\Bigl{(}0.14\kappa-2.12\gamma^{2}-3.35\Bigr{)}\varphi(1.96)+O(n^{-2}),$ (3)
where $\varphi$ is the $\mathcal{N}(0,1)$ probability density function, and the constants have been rounded. The row labeled ‘Norm’ in Table 1 of Hall, (1988) contains $3.350.14\kappa-2.12\gamma^{2}-$ 3.35, omitting the factor $2\varphi(1.96)/n$ common to all of the coverage error expressions.
Table 1 is missing a row for the usual Student’s $t$ interval in equation (1). That interval differs from the one for (2) in using $s$ instead of $\hat{\sigma}$ and in using $t^{1-\alpha}_{(n-1)}$ instead of $\Phi^{-1}(1-\alpha)$.
Section 2 rederives Hall’s formula for coverage errors in 95% confidence intervals based on the Gaussian interval. Section 3 then constructs formulas for the Student $t$ interval. The key step is to use an asymptotic formula from Abramowitz and Stegun, (1972) relating quantiles of Student’s $t$ distribution to those of the $\mathcal{N}(0,1)$ distribution. Both tasks were made simpler by noticing that Hall’s derivation for the ‘Norm’ interval does not use all of the machinery that the bootstrap intervals in that paper require.
2 Derivation
Here we look at the derivation for the row called ‘Norm’ before constructing the analogous formula for the standard $t$ intervals. Let $z_{\alpha}=\Phi^{-1}(\alpha)$ for $10<\alpha<$ 1 and then let $C(\alpha)=\Pr(\mu\leqslant\hat{\theta}+n^{-1/2}\hat{\sigma}z_{\alpha})$ be the one sided coverage probability for the normal interval $(-\infty,\hat{\theta}+n^{-1/2}\hat{\sigma}z_{\alpha}]$. From page 949 of Hall, (1988),
$C(\alpha)=\alpha-n^{-1/2}\frac{1}{6}\gamma(2z_{\alpha}^{2}+1)\varphi(z_{\alpha})+n^{-1}z_{\alpha}\Bigl{(}\frac{1}{12}\kappa(z_{\alpha}^{2}-3)\\ -\frac{1}{18}\gamma^{2}\bigl{(}z_{\alpha}^{4}+2z_{\alpha}^{2}-3\bigr{)}-\frac{1}{4}\bigl{(}z_{\alpha}^{2}+3\bigr{)}\Bigr{)}\varphi(z_{\alpha})+O(n^{-3/2}).$ (4)
For $\alpha<1/2$, the two-sided coverage for
$[\hat{\theta}-n^{-1/2}\hat{\sigma}z_{1-\alpha},\hat{\theta}+n^{-1/2}\hat{\sigma}z_{1-\alpha}]$
with $z_{1-\alpha}=-z_{\alpha}$ is $C(1-\alpha)-C(\alpha)$ which equals
$1-2\alpha+2n^{-1}z_{1-\alpha}\Bigl{(}\frac{1}{12}\kappa(z_{1-\alpha}^{2}-3)\\ -\frac{1}{18}\gamma^{2}\bigl{(}z_{1-\alpha}^{4}+2z_{1-\alpha}^{2}-3\bigr{)}-\frac{1}{4}\bigl{(}z_{1-\alpha}^{2}+3\bigr{)}\Bigr{)}\varphi(z_{1-\alpha})+O(n^{-3/2}).$
For the bootstrap methods covered by (3) the $O(n^{-3/2})$ error is actually $O(n^{-2})$ because the $O(n^{-3/2})$ terms cancel the same way that the $O(n^{-1/2})$ terms do. This may very well be true for ‘Norm’ as well, but here we only consider terms up to order $O(1/n)$.
On page 949, Hall, (1988) writes: “The term $-3.35$ appearing in the third column of Table 1 arises from the difference between the standard normal distribution function and an expansion of Student’s t distribution function.” This suggests that all we must do is drop the intercept from the expression for the normal theory result, but for the avoidance of doubt we work through the expansion here to verify that this is the case.
The coverage error is the attained coverage minus the nominal coverage. For the two sided interval the coverage error is then
$\frac{2z_{1-\alpha}}{n}\Bigl{(}\frac{z_{\alpha}^{2}-3}{12}\kappa-\frac{z_{\alpha}^{4}+2z_{\alpha}^{2}-3}{18}\gamma^{2}-\frac{z_{\alpha}^{2}+3}{4}\Bigr{)}\varphi(z_{\alpha})+O(n^{-3/2})$ (5)
where even powers of $z_{1-\alpha}$ and $z_{\alpha}$ are equal because $z_{\alpha}=-z_{1-\alpha}$. The corresponding entry for Table 1, which leaves out the factor $2\varphi(z_{\alpha})/n$ is then
$z_{1-\alpha}\frac{z_{\alpha}^{2}-3}{12}\kappa-z_{1-\alpha}\frac{z_{\alpha}^{4}+2z_{\alpha}^{2}-3}{18}\gamma^{2}-z_{1-\alpha}\frac{z_{\alpha}^{2}+3}{2}.$ (6)
Using $z_{1-\alpha}=z_{0.975}\doteq 1.96$ we get $3.350.14\kappa-2.12\gamma^{2}-$ 3.35 after rounding, matching Hall’s Table 1.
For Gaussian data, we have $\gamma=\kappa=0$. Then the standard normal theory intervals (2) have coverage error
$-\frac{2\varphi(z_{\alpha})}{n}z_{1-\alpha}\frac{z_{\alpha}^{2}+3}{2}+O(n^{-3/2})\doteq-\frac{0.78}{n}.$
The negative sign indicates undercoverage. The Student $t$ intervals (1) have coverage error zero. Therefore we anticipate, and we will see, that the counterpart to (6) for Student $t$ intervals equals zero when $\gamma=\kappa=0$.
3 Adjusted interval
To get the required formula, we need to replace $z_{1-\alpha}$ by $t_{(n-1)}^{1-\alpha}$, the $1-\alpha$ quantile of Student’s $t$ distribution. We also have to replace $\hat{\sigma}$ by $s$. We place $1-\alpha$ in the superscript of $t^{1-\alpha}_{(n-1)}$ so it does not collide with the degrees of freedom $n-1$, while placing $1-\alpha$ in ths subscript of $z_{1-\alpha}$ so that it will not collide with exponents when we raise that quantile to integer powers.
Page 949 of Abramowitz and Stegun, (1972) gives an asymptotic expansion relating the $t$ quantiles to the Gaussian ones:
$t_{(\nu)}^{\alpha}=z_{\alpha}+\frac{g_{1}(z_{\alpha})}{\nu}+\frac{g_{2}(z_{\alpha})}{\nu^{2}}+\frac{g_{3}(z_{\alpha})}{\nu^{3}}+\frac{g_{4}(z_{\alpha})}{\nu^{4}}+\cdots$ (7)
where
$g_{1}(x)$ $=\frac{x^{3}+x}{4},$
$g_{2}(x)$ $=\frac{5x^{5}+16x^{3}+3x}{96},$
$g_{3}(x)$ $=\frac{3x^{7}+19x^{5}+17x^{3}-15x}{384},\quad\text{and}$
$g_{4}(x)$ $=\frac{79x^{9}+776x^{7}+1482x^{5}-1920x^{3}-945x}{92160}.$
Using the first two terms of expansion (7),
$t^{\alpha}_{(n-1)}=z_{\alpha}+\frac{z_{\alpha}^{3}+z_{\alpha}}{4(n-1)}+O(n^{-2})=z_{\alpha}+\frac{z_{\alpha}^{3}+z_{\alpha}}{4n}+O(n^{-2}).$
Similarly
$s=\hat{\sigma}\sqrt{\frac{n}{n-1}}=1+\frac{1}{2n}+O(n^{-2}),$
and so
$st^{\alpha}_{(n-1)}$ $=\hat{\sigma}z_{\alpha}\Bigl{(}1+\frac{1}{2n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}\Bigr{)}\Bigl{(}1+\frac{z_{\alpha}^{2}+1}{4n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}\Bigr{)}$
$=\hat{\sigma}z_{\alpha}\Bigl{(}1+\frac{z_{\alpha}^{2}+3}{4n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}\Bigr{)}$
$=\hat{\sigma}z_{\alpha^{\prime}}$
where
$\alpha^{\prime}$ $=\Phi(z_{\alpha^{\prime}})=\Phi\Bigl{(}z_{\alpha}+\frac{z_{\alpha}^{3}+3z_{\alpha}}{4n}+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}\Bigr{)}$
$=\alpha+\frac{z_{\alpha}^{3}+3z_{\alpha}}{4n}\varphi(z_{\alpha})+O\Bigl{(}\frac{1}{n^{2}}\Bigr{)}.$
Using (5), the coverage error of the two sided Student $t$ interval is
$\phantom{=}\ 2(\alpha-\alpha^{\prime})+\frac{z_{1-\alpha^{\prime}}}{n}\Bigl{(}\frac{z_{\alpha^{\prime}}^{2}-3}{6}\kappa-\frac{z_{\alpha^{\prime}}^{4}+2z_{\alpha^{\prime}}^{2}-3}{9}\gamma^{2}-\frac{z_{\alpha^{\prime}}^{2}+3}{2}\Bigr{)}\varphi(z_{\alpha^{\prime}})+O(n^{-3/2})$
$=2(\alpha-\alpha^{\prime})+\frac{z_{1-\alpha}}{n}\Bigl{(}\frac{z_{\alpha}^{2}-3}{6}\kappa-\frac{z_{\alpha}^{4}+2z_{\alpha}^{2}-3}{9}\gamma^{2}-\frac{z_{\alpha}^{2}+3}{2}\Bigr{)}\varphi(z_{\alpha})+O(n^{-3/2}).$
Now
$2(\alpha-\alpha^{\prime})$ $=-\frac{z_{\alpha}^{3}+3z_{\alpha}}{2n}+O(n^{-2})=\frac{z_{1-\alpha}}{n}\Bigl{(}\frac{z_{\alpha}^{2}+3}{2n}\Bigr{)}+O(n^{-2}).$
It follows that the coverage error in Student’s $t$ interval is
$\frac{z_{1-\alpha}}{n}\Bigl{(}\frac{z_{\alpha}^{2}-3}{6}\kappa-\frac{z_{\alpha}^{4}+2z_{\alpha}^{2}-3}{9}\gamma^{2}\Bigr{)}\varphi(z_{\alpha})+O(n^{-3/2}).$
The coefficients of $\kappa$ and $\gamma^{2}$ matches those of equation (5) for the normal theory method and the intercept term is zero.
Acknowledgments
This work was supported by the U.S. National Science Foundation under grant DMS-2152780.